# TowerVision: Understanding and Improving Multilinguality in Vision-Language Models

### v2

@@ 3.2 Baselines
For evaluation, we leverage the lmms-eval framework (Zhang et al., 2025b), which enables a systematic comparison of TowerVision against leading open VLMs. We include several multilingual multimodal models, such as CulturalPangea-7B (Yue et al., 2025), designed to address gaps in multilingual cultural understanding, and Aya-Vision-8B (Singh et al., 2024), optimized for a broad range of vision-language tasks. In addition, we evaluate models from the Gemma3-Instruct (Gemma3-4B-it, Gemma3-12B-it; Team et al. 2025) and the Qwen2.5-VL-Instruct families (Qwen2.5-VL-3B-Instruct, Qwen2.5-VL-7B-Instruct; Qwen et al. 2025), both of which have demonstrated strong performance across a variety of multimodal benchmarks. Finally, we report results for a LLaVA-based model, Llama3-Llava-Next-8B (Liu et al., 2024a), a general-purpose VLM with strong performance across a wide range of tasks. The exact checkpoints for all models are listed in §A.2.
For TowerVideo, we consider several competitive open-source video models of comparable scale, including VideoLLaMA3-7B (Zhang et al., 2025a), LLaVA-Video-7B (Zhang et al., 2025c) also trained on LLaVA-Video-178k, and ViMUL-7B (Shafique et al., 2025), a multilingual video model.
@@
Table 2: Multimodal Translation Benchmarks. We report xComet (Guerreiro et al., 2024) for Multi30K and contrastive pairwise accuracy for CoMMuTE.
@@ 3.3 Main Results
TowerVision is less competitive on OCR-related tasks. We hypothesize this is likely due to the limited amount of OCR-focused data in VisionBlocks compared against other models. Since we primarily pretrained TowerVision on large-scale image-caption datasets emphasizing natural images and language alignment, it struggles with scanned text or OCR-heavy scenarios. Despite these limitations, TowerVision does obtain superior performance compared against AyaVision-8B and Llam3LLaVANext-8B, the former of which has seen significant amounts of OCR-specific data (Singh et al., 2024).
@@
Multilingual fine-tuning improves cross-lingual performance in TowerVideo. In Table 3, we report averages across multiple-choice accuracy and open-ended responses, which are automatically judged using GPT-4o (OpenAI et al., 2024), with the same evaluation prompt as Shafique et al. (2025). We compare our TowerVideo models, including the 9B variant, to strong open-source baselines. Our multilingual models are competitive across several languages despite using smaller datasets and fewer frames (for instance, VideoLLaMA3 uses 180 frames). Specifically, ViMUL was trained with separate copies of the dataset for each language, whereas our approach uses a single copy with half in English and the other half uniformly translated into the supported languages. Overall, these results highlight the effectiveness of video-based multilingual fine-tuning in improving cross-lingual reasoning.
@@ 4 Where and How Does Multilinguality Matter?
Multilingual backbones improve cross-modal performance. The choice of backbone in TowerVision can substantially influence performance across multilingual and multimodal tasks. We focus on two complementary aspects. First, we examine the significance of multilingual capacity by comparing the Tower+ backbone, which is highly multilingual and designed for general-purpose multilingual text tasks, against Gemma2, the model on which Tower+ was built. Second, we investigate the impact of instruction tuning before modality fusion, which is widely applied in modern VLMs from the start (Liu et al., 2023b; Bai et al., 2025a), but whose effect on the final model remains unclear. To study these effects, we train TowerVision at 2B and 9B scales using four backbones: Gemma2-pt (pretrained, not instruction-tuned), Gemma2-it (instruction-tuned), Tower+pt (pretrained Tower+), and Tower+it (instruction-tuned Tower+), following the recipe in §2. As shown in Table 4, using Tower+ consistently outperforms Gemma2, confirming the importance of a strong multilingual backbone for robust cross-modal understanding. At smaller scales, non-instructed models (Gemma2-pt, Tower+pt) retain stronger raw visual extraction, while instruction-tuned variants excel in cultural knowledge and reasoning. By the 9B scale, this gap narrows, with instruction-tuned models integrating both skills and achieving state-of-the-art performance. These findings underscore the complementary roles of multilingual pretraining and instruction tuning, and the need for careful backbone selection in VLMs.
@@
Expanding languages improves cross-lingual generalization in VLMs. We study how language coverage in training data impacts performance on both included and excluded languages. Specifically, we compare training on 10 high-resource “core languages” versus the full set of languages, while controlling for dataset size. Our questions are: (i) whether adding balanced multimodal data for more languages improves performance on core languages (Conneau et al., 2020; Hu et al., 2020), and (ii) whether unsupported languages benefit in zero-shot fashion if related languages are present (Ni et al., 2021). We train TowerVision at 2B and 9B scales using the recipe in §2, first on 10 “core” languages (English, German, Dutch, Portuguese, Russian, Simplified and Traditional Chinese, Spanish, French, Italian), then on all available languages. Results in Figure 3 (more details in §A.4) show that broader language coverage consistently improves performance, with larger gains at the 2B scale. Zero-shot improvements for unsupported languages further support cross-lingual transfer when related languages are included. These findings highlight the value of expanding multilingual data, particularly for smaller models.

### v3

@@ 3.2 Baselines
For evaluation, we leverage the lmms-eval framework (lmms_eval2024), which enables a systematic comparison of TowerVision against leading open VLMs. We include several multilingual multimodal models, such as CulturalPangea-7B (yue2025pangeafullyopenmultilingual), designed to address gaps in multilingual cultural understanding, and Aya-Vision-8B (singh2024aya), optimized for a broad range of vision-language tasks. In addition, we evaluate models from the Gemma3-Instruct (Gemma3-4B-it, Gemma3-12B-it; team2025gemma) and the Qwen2.5-VL-Instruct families (Qwen2.5-VL-3B-Instruct, Qwen2.5-VL-7B-Instruct; qwen2025qwen25technicalreport), both of which have demonstrated strong performance across a variety of multimodal benchmarks. Finally, we report results for a LLaVA-based model, Llama3-Llava-Next-8B (liu2024llavanext), a general-purpose VLM with strong performance across a wide range of tasks. The exact checkpoints for all models are listed in §A.2.
For TowerVideo, we consider several competitive open-source video models of comparable scale, including VideoLLaMA3-7B (damonlpsg2025videollama3), LLaVA-Video-7B (zhang2025llavavideovideoinstructiontuning) also trained on LLaVA-Video-178k, and ViMUL-7B (shafique2025culturallydiversemultilingualmultimodalvideo), a multilingual video model.
@@
Table 2: Multimodal Translation Benchmarks. We report xComet (guerreiro2024xcomet) for Multi30K and contrastive pairwise accuracy for CoMMuTE.
@@ 3.3 Main Results
TowerVision is less competitive on OCR-related tasks. We hypothesize this is likely due to the limited amount of OCR-focused data in VisionBlocks compared against other models. Since we primarily pretrained TowerVision on large-scale image-caption datasets emphasizing natural images and language alignment, it struggles with scanned text or OCR-heavy scenarios. Despite these limitations, TowerVision does obtain superior performance compared against AyaVision-8B and Llam3LLaVANext-8B, the former of which has seen significant amounts of OCR-specific data (singh2024aya).
@@
Multilingual fine-tuning improves cross-lingual performance in TowerVideo. In Table 3, we report averages across multiple-choice accuracy and open-ended responses, which are automatically judged using GPT-4o (openai2024gpt4ocard), with the same evaluation prompt as shafique2025culturallydiversemultilingualmultimodalvideo. We compare our TowerVideo models, including the 9B variant, to strong open-source baselines. Our multilingual models are competitive across several languages despite using smaller datasets and fewer frames (for instance, VideoLLaMA3 uses 180 frames). Specifically, ViMUL was trained with separate copies of the dataset for each language, whereas our approach uses a single copy with half in English and the other half uniformly translated into the supported languages. Overall, these results highlight the effectiveness of video-based multilingual fine-tuning in improving cross-lingual reasoning.
@@ 4 Where and How Does Multilinguality Matter?
Multilingual backbones improve cross-modal performance. The choice of backbone in TowerVision can substantially influence performance across multilingual and multimodal tasks. We focus on two complementary aspects. First, we examine the significance of multilingual capacity by comparing the Tower+ backbone, which is highly multilingual and designed for general-purpose multilingual text tasks, against Gemma2, the model on which Tower+ was built. Second, we investigate the impact of instruction tuning before modality fusion, which is widely applied in modern VLMs from the start (liu2023visualinstructiontuning; bai2025qwen25vltechnicalreport), but whose effect on the final model remains unclear. To study these effects, we train TowerVision at 2B and 9B scales using four backbones: Gemma2-pt (pretrained, not instruction-tuned), Gemma2-it (instruction-tuned), Tower+pt (pretrained Tower+), and Tower+it (instruction-tuned Tower+), following the recipe in §2. As shown in Table 4, using Tower+ consistently outperforms Gemma2, confirming the importance of a strong multilingual backbone for robust cross-modal understanding. At smaller scales, non-instructed models (Gemma2-pt, Tower+pt) retain stronger raw visual extraction, while instruction-tuned variants excel in cultural knowledge and reasoning. By the 9B scale, this gap narrows, with instruction-tuned models integrating both skills and achieving state-of-the-art performance. These findings underscore the complementary roles of multilingual pretraining and instruction tuning, and the need for careful backbone selection in VLMs.
@@
Expanding languages improves cross-lingual generalization in VLMs. We study how language coverage in training data impacts performance on both included and excluded languages. Specifically, we compare training on 10 high-resource “core languages” versus the full set of languages, while controlling for dataset size. Our questions are: (i) whether adding balanced multimodal data for more languages improves performance on core languages (conneau2020unsupervisedcrosslingualrepresentationlearning; hu2020xtrememassivelymultilingualmultitask), and (ii) whether unsupported languages benefit in zero-shot fashion if related languages are present (ni2021m3plearninguniversalrepresentations). We train TowerVision at 2B and 9B scales using the recipe in §2, first on 10 “core” languages (English, German, Dutch, Portuguese, Russian, Simplified and Traditional Chinese, Spanish, French, Italian), then on all available languages. Results in Figure 3 (more details in §A.4) show that broader language coverage consistently improves performance, with larger gains at the 2B scale. Zero-shot improvements for unsupported languages further support cross-lingual transfer when related languages are included. These findings highlight the value of expanding multilingual data, particularly for smaller models.